# Pair-density-wave superconductivity and Anderson’s theorem in bilayer nickelates

### v1

@@ Abstract
Fig. 3(c) demonstrates that the transition between the uniform SC and PDW phases is first-order, while the transitions between the PDW and normal phases, or between the SC and normal phases, are continuous. Although the form factor in the interlayer hopping $t_{\perp}$ is motivated by the specific band structure of bilayer nickelates, we find that it does not qualitatively alter the phase diagram. Consequently, similar PDW phases could be realized more generally in other bilayer systems subject to an external electric field. Finally, Fig. 4 examines the dependence on electron filling, demonstrating that the PDW phase remains robust against doping.
@@
where $T_{c}$ is linearly suppressed by the effective scattering rate $\Gamma=n_{\mathrm{imp}}V^{2}$, and $T_{c,0}$ denotes the critical temperature in the absence of disorder. At $t_{\perp}=0$, the dimensionless depairing coefficient $\alpha_{\mathrm{dis}}$ is given by,
$\alpha_{\mathrm{dis}}$ $=$ $\left\langle\mathrm{Tr}\left(\tilde{F}_{c}^{\dagger}(k)\tilde{F}_{c}(k)\right)\right\rangle_{\mathrm{FS}},$ (10)
@@
$F_{c}$ $=$ $V_{\mathrm{dis}}\Delta-\Delta V_{\mathrm{dis}}^{*}.$ (11)
Here, the tilde denotes normalization by the disorder strength and the superconducting gap. The depairing coefficient $\alpha_{\mathrm{dis}}$ quantifies the robustness of the superconducting state, where a larger $\alpha_{\mathrm{dis}}$ corresponds to a more fragile state, while $\alpha_{\mathrm{dis}}=0$ indicates complete robustness in the weak-disorder limit.
Disorder $\mathcal{M}$ $\mathcal{T}$ $t_{\perp}=0$ $t_{\perp}\neq 0$
@@
Table 1: Summary of the generalized Anderson theorem for interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$. The second and third columns list the symmetry properties under mirror reflection $\mathcal{M}$ and time-reversal symmetry $\mathcal{T}$ for the four disorder potentials. The fourth and fifth columns indicate the stability of interlayer pairing for the cases without ($t_{\perp}=0$) and with ($t_{\perp}\neq 0$) interlayer hopping, determined by the depairing coefficient $\alpha_{\mathrm{dis}}$. A ✓  denotes $\alpha_{\mathrm{dis}}=0$ (stable SC), while a ✗  denotes $\alpha_{\mathrm{dis}}\neq 0$ (fragile SC).
For $t_{\perp}=0$ and interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$, we derive that $F_{c}=0$ when the disorder potential possesses $\mathcal{MT}$ symmetry. This implies that interlayer spin-singlet superconductivity is stable under $\mathcal{M}\mathcal{T}$-preserving disorder scattering. Accordingly, Anderson’s theorem for interlayer pairing should be modified to require $\mathcal{M}\mathcal{T}$ symmetry preservation. Applying the same analysis to intralayer pairing, $\Delta_{\parallel}=\tau_{0}\otimes i\sigma_{y}$, we recover the conventional Anderson theorem, where the non-pair-breaking condition is simply the preservation of $\mathcal{T}$.
Another important question is how this generalized Anderson theorem is modified in the presence of finite interlayer hopping $t_{\perp}$. To address this, we further analyze $\alpha_{\mathrm{dis}}$ for $t_{\perp}\neq 0$. In this case, $\alpha_{\mathrm{dis}}$ is no longer determined solely by the commutation relation between the pairing gap and the disorder potential. Instead, additional contributions arise from the commutator between the normal-state Green’s function and the pairing gap. The results are summarized in Table 1.
As shown, only the $V_{2}$ disorder channel, which is time-reversal odd and corresponds to current-like disorder, is affected by finite $t_{\perp}$ and becomes pair-breaking. Consequently, the generalized Anderson theorem derived here remains applicable in realistic experimental settings, where disorder is typically time-reversal symmetric. Notably, for mirror-symmetry breaking disorder ($V_{3}$), $\alpha_{\mathrm{dis}}$ is nonzero for both $t_{\perp}=0$ and $t_{\perp}\neq 0$, though its quantitative value changes due to the non-zero commutator with the normal Green’s function. In real materials, individual defects likely break mirror symmetry; therefore, $T_{c}$ in bilayer nickelates is expected to be sensitive to disorder, and even higher $T_{c}$ values may be achievable in cleaner samples.
@@
$H_{\mathrm{eff}}$ $=$ $\sum_{\bm{k},\sigma}\phi_{\bm{k},\sigma}^{\dagger}\left(\begin{array}[]{cc}\xi(\bm{k})+D&-\gamma(\bm{k})\\ \gamma_{(}\bm{k})&\xi(\bm{k})-D\end{array}\right)\phi_{\bm{k},\sigma}=\sum_{\bm{k}}\phi_{\bm{k},\sigma}^{\dagger}\mathcal{H}_{N}(\bm{k})\phi_{\bm{k},\sigma}$ (S4)
@@
We also present the zero-temperature mean-field phase diagram without the momentum-dependent form factor, taking $t_{\perp}(\bm{k})=t_{\perp}$. The resulting phase diagram exhibits the same qualitative features as that obtained with the form factor shown in Fig. 3(a). As illustrated in Fig. S1, the results show that PDW is more stable than without form factor case. This is related to the shape of Fermi-surface, where the nesting vector is more well-defined in the absence of the form factor.
@@
In this section, we provide derivation of Eq. S16 and expression of depairing coefficient $\alpha_{\mathrm{dis}}$ based on the 1st order Born-approximation. $\alpha_{\mathrm{dis}}$ quantifies the slope of suppression of the critical temperature under disorder,
$\log\left(\frac{T_{c}}{T_{c,0}}\right)=-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (S16)
where $T_{c}$ is linearly suppressed by the effective scattering rate $\Gamma=n_{\mathrm{imp}}V^{2}$. Here $T_{c,0}$ denotes the critical temperature in the absence of disorder.
C.1 A. Setup and 1st order Born approximation
@@
where $\tilde{V}_{\mathrm{dis}}$ denotes the disorder potential written in the Nambu basis, and $i\omega_{n}$ is the fermionic Matsubara frequency.
@@
where $G$ and $F$ denote the diagonal and off-diagonal blocks of the BdG Green’s function $\mathcal{G}_{\mathrm{BdG}}$, respectively. The limit $\Delta_{0}\rightarrow 0$ is taken since we focus on the critical temperature, assuming that the superconducting transition is a continuous second-order phase transition.
C.2 B. $t_{\perp}=0$ case
We start with the simple case $t_{\perp}=0$. In this limit, the two bands are degenerate with dispersion $\xi_{\bm{k}}$. The normal and anomalous Green’s functions $G$ and $F$ in the limit $\Delta_{0}\rightarrow 0$ are simplified as
@@
Substituting these expressions into Eqs. S24 and S25, we obtain
$\Sigma_{\omega}$ $=$ $-\frac{i\omega_{n}}{|\omega_{n}|}\,\Gamma,\quad\Sigma_{\Delta}=\frac{\Delta_{0}}{|\omega_{n}|}\Gamma\left(1-\frac{\alpha_{\mathrm{dis}}}{4}\right),$
which lead to the renormalized frequency and pairing amplitude,
$\tilde{\omega}_{n}$ $=$ $\omega_{n}+\frac{\tilde{\omega}_{n}}{|\tilde{\omega}_{n}|}\Gamma,\quad\tilde{\Delta}_{0}=\Delta_{0}+\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|}\Gamma\left(1-\frac{\alpha_{\mathrm{dis}}}{4}\right).$ (S27)
Here $\alpha_{\mathrm{dis}}$ denotes the depairing coefficient, and $F_{c}$ is the superconducting fitness function defined as
$\alpha_{\mathrm{dis}}$ $=$ $\left\langle\mathrm{Tr}\left(\tilde{F}_{c}^{\dagger}(k)\tilde{F}_{c}(k)\right)\right\rangle_{\mathrm{FS}},$ (S28)
@@
In the above derivation, we have used the identity
[…]
Considering the disorder channels ${\widehat{V}}=\tau_{i}$ defined in Eq. S21 and the interlayer pairing ${\widehat{\Delta}}=\tau_{1}$, we find that
@@
This leads directly to the results summarized in Table 1. Note that the $\tau_{3}$ disorder channel breaks the composite $\mathcal{M}\mathcal{T}$ symmetry, while all other channels preserve it.
C.2.1 Renormalized gap equation and $T_{c}$
@@
where $g>0$ denotes the attractive pairing interaction.
Replacing the momentum summation by an integral over energies near the Fermi surface, the gap equation can be rewritten as
$\frac{1}{2g}\simeq\pi N(0)T_{c,0}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{1}{|\omega_{n}|}\;\rightarrow\;\frac{\Delta_{0}}{2g}\simeq\pi N(0)T_{c}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|},$
where $N(0)$ is the density of states at the Fermi level and $\Lambda_{UV}$ is an ultraviolet energy cutoff imposed on the Matsubara frequencies. The angular average over the Fermi surface is defined as $\langle\mathcal{O}\rangle_{\Omega}=\int_{FS}\frac{d\Omega}{4\pi}\mathcal{O}$.
Substituting the renormalized parameters $(\tilde{\omega}_{n},\tilde{\Delta}_{0})$, we obtain
$\frac{1}{2gN(0)}\simeq 2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\Bigg[\frac{1}{\omega_{n}}-\frac{\Gamma}{\omega_{n}^{2}}\alpha_{\mathrm{dis}}\Bigg],$ (S33)
where $\Gamma=n_{\mathrm{imp}}V^{2}$ is the effective scattering rate.
To perform the Matsubara summation in Eq. S33, we make use of the digamma function $\Psi(z)$ and its derivative $\Psi^{(1)}(z)\equiv d\Psi(z)/dz$, defined as
@@
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}}$ $=$ $-\Bigg[\Psi\!\left(\frac{1}{2}\right)-\Psi\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$
@@
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}^{2}}$ $=$ $\frac{1}{2\pi T_{c}}\Bigg[\Psi^{(1)}\!\left(\frac{1}{2}\right)-\Psi^{(1)}\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$
@@
$\log\left(\frac{T_{c}}{T_{c,0}}\right)=-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (S34)
@@
C.3 C. $t_{\perp}\neq 0$ case
For the case $t_{\perp}\neq 0$, the form of Eq. S26 is modified due to the band splitting, $\xi_{\pm}(\bm{k})=\xi(\bm{k})\pm|t_{\perp}|$. As a result, the final expression for the depairing coefficient $\alpha_{\mathrm{dis}}$ is also modified.
Following the setup introduced in Sec. C.1, we consider the normal and anomalous Green’s functions $G$ and $F$ in the limit $\Delta_{0}\rightarrow 0$. The normal Green’s function can be written as
$\lim_{\Delta_{0}\rightarrow 0}G(\bm{k},i\omega_{n})$ $=$ $(i\omega_{n}-\xi_{+})^{-1}\tilde{\tau}_{0}+(i\omega_{n}-\xi_{-})^{-1}\tilde{\tau}_{3}\equiv G_{N}(\bm{k},i\omega_{n}),$ (S35)
which is expressed in the band basis. Here we use the tilde notation for Pauli matrices to emphasize that they are defined in the band basis, as introduced in Eq. S22.
@@
The anomalous Green’s function in the same limit is given by
$\lim_{\Delta_{0}\rightarrow 0}\frac{F(\bm{k},i\omega_{n})}{\Delta_{0}}$ $=$ $-G_{N}{\widehat{\Delta}}G_{N}^{*}=-G_{N}G_{N}^{*}{\widehat{\Delta}}$, (S36)
@@
where we have used the fact that $G_{N},{\widehat{\Delta}}]=0$, since both operators are proportional to ${\widehat{\Delta}}=\tilde{\tau}_{3}$.
As shown in the previous section, the key quantity governing the robustness of superconductivity within Anderson’s theorem is the relative renormalization of the frequency and pairing self-energies. In particular, when the two renormalizations are identical (i.e., $\alpha_{\mathrm{dis}}=0$), disorder does not suppress the critical temperature $T_{c}$.
We now evaluate Eqs. S24 and S25 for the case $t_{\perp}\neq 0$, again considering interlayer pairing with ${\widehat{\Delta}}=\tilde{\tau}_{3}$. We find
$\lim_{\Delta_{0}\rightarrow 0}\Big[{\widehat{V}}\,G(\bm{k},i\omega_{n})\,{\widehat{V}}\Big$] $=$ $\Big[{\widehat{V}}\,G_{N}\,{\widehat{V}}\Big$] (S37)
$=$ $\Big[G_{N}{\widehat{V}}+[{\widehat{V}},G_{N}]\Big]{\widehat{V}}$,
and

### v2

@@ Abstract
Fig. 3(c) demonstrates that the transition between the uniform SC and PDW phases is first-order, while the transitions between the PDW and normal phases, or between the SC and normal phases, are continuous. Although the form factor in the interlayer hopping $t_{\perp}$ is motivated by the specific band structure of bilayer nickelates, we find that it does not qualitatively alter the phase diagram and PDW is more stabilized in case of without form factore, due to better nesting condition (See Fig. S1). Consequently, similar PDW phases could be realized more generally in other bilayer systems subject to an external electric field. Finally, Fig. 4 examines the dependence on electron filling, demonstrating that the PDW phase remains robust against doping.
@@
where $T_{c}$ is linearly suppressed by the effective scattering frequency, $\Gamma\equiv\pi N(0)n_{\mathrm{imp}}V^{2}$ is used. $T_{c,0}$ denotes the critical temperature in the absence of disorder and $N(0)$ is the density of states at the Fermi energy. The dimensionless depairing coefficient $\alpha_{\mathrm{dis}}$ is given by,
$\alpha_{\mathrm{dis}}$ $=$ $\frac{1}{4}\mathrm{Tr}\left(\tilde{F}_{c}^{\dagger}\ \tilde{F}_{c}\right),$ (10)
@@
$\tilde{F}_{c}$ $=$ ${\widehat{V}}_{\mathrm{dis}}{\widehat{\Delta}}-{\widehat{\Delta}}{\widehat{V}}_{\mathrm{dis}}^{*}.$ (11)
It is important to not that ${\widehat{V}}_{\mathrm{dis}}$, $^{\widehat{\Delta}}$ denotes normalized quantity divided by the disorder strength and the superconducting gap. The depairing coefficient $\alpha_{\mathrm{dis}}$ quantifies the robustness of the superconducting state, where a larger $\alpha_{\mathrm{dis}}$ corresponds to a more fragile state, while $\alpha_{\mathrm{dis}}=0$ indicates complete robustness in the weak-disorder limit.
Disorder $\mathcal{M}$ $\mathcal{T}$ Stab. of $\Delta_{\perp}$
@@
Table 1: Summary of the generalized Anderson theorem for interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$. The second and third columns list the symmetry properties under mirror reflection $\mathcal{M}$ and time-reversal symmetry $\mathcal{T}$ for the four disorder potentials. The fourth column indicate the stability of interlayer pairing for both cases without ($t_{\perp}=0$) and with ($t_{\perp}\neq 0$) interlayer hopping, determined by the depairing coefficient $\alpha_{\mathrm{dis}}$. A ✓  denotes $\alpha_{\mathrm{dis}}=0$ (stable SC), while a ✗  denotes $\alpha_{\mathrm{dis}}\neq 0$ (fragile SC).
For $t_{\perp}=0$ and interlayer spin-singlet pairing, $\Delta_{\perp}=\tau_{x}\otimes i\sigma_{y}$, we derive that $F_{c}=0$ when the disorder potential possesses $\mathcal{MT}$ symmetry. This implies that interlayer spin-singlet superconductivity is stable under $\mathcal{M}\mathcal{T}$-preserving disorder scattering. Accordingly, Anderson’s theorem for interlayer pairing should be modified to require $\mathcal{M}\mathcal{T}$ symmetry preservation. Applying the same analysis to intralayer pairing, $\Delta_{\parallel}=\tau_{0}\otimes i\sigma_{y}$, we recover the conventional Anderson theorem, where the non-pair-breaking condition is simply the preservation of $\mathcal{T}$. The results are summarized in Table 1.
Another important question is how this generalized Anderson theorem is modified in the presence of finite interlayer hopping $t_{\perp}$. To address this, we extend our analysis of $\alpha_{\mathrm{dis}}$ to the $t_{\perp}\neq 0$ case. Our derivation reveals that the expression for $\alpha_{\mathrm{dis}}$ (Eq. 10) remains invariant even under the band splitting induced by $t_{\perp}$. Hence underlying symmetry plays a more important role in determining robustness under weak disorder than the specific microscopic details of the Hamiltonian.
Consequently, the generalized Anderson theorem established here remains applicable in realistic experimental settings where disorder is typically time-reversal symmetric. Notably, for mirror-symmetry-breaking disorder ($V_{3}$), $\alpha_{\mathrm{dis}}$ remains non-zero for both $t_{\perp}=0$ and $t_{\perp}\neq 0$, even though the clean-limit critical temperature $T_{c,0}$ is rescaled by the modified density of states (DOS). In real materials, individual defects are likely to break mirror symmetry locally. Therefore, $T_{c}$ in bilayer nickelates is expected to be sensitive to disorder, and higher $T_{c}$ values can be achieved in cleaner samples.
@@
$H_{\mathrm{eff}}$ $=$ $\sum_{\bm{k},\sigma}\phi_{\bm{k},\sigma}^{\dagger}\left(\begin{array}[]{cc}\xi(\bm{k})+D&\gamma(\bm{k})\\ \gamma_{(}\bm{k})&\xi(\bm{k})-D\end{array}\right)\phi_{\bm{k},\sigma}=\sum_{\bm{k}}\phi_{\bm{k},\sigma}^{\dagger}\mathcal{H}_{N}(\bm{k})\phi_{\bm{k},\sigma}$ (S4)
@@
We also present the zero-temperature mean-field phase diagram without the momentum-dependent form factor, taking $\gamma(\bm{k})=t_{\perp}$. The resulting phase diagram exhibits the same qualitative features as that obtained with the form factor shown in Fig. 3(a). As illustrated in Fig. S1, the results show that PDW is more stable than without form factor case. This is related to the shape of Fermi-surface, where the nesting vector is more well-defined in the absence of the form factor.
@@
C.1 A. Setup
@@
where $G_{\mathrm{BdG}}(\bm{k},i\omega_{n})=(i\omega_{n}-H_{\mathrm{BdG}})^{-1}$ is BdG Green’s function and $\tilde{V}_{\mathrm{dis}}$ denotes the disorder potential written in the Nambu basis. $i\omega_{n}$ is the fermionic Matsubara frequency. A diagrammatic representation is shown in Fig. S1(a).
@@
where $G$ and $F$ denote the diagonal and off-diagonal blocks of the BdG Green’s function $\mathcal{G}_{\mathrm{BdG}}$, respectively. The limit $\Delta_{0}\rightarrow 0$ is taken since we will focus on the critical temperature, assuming that the superconducting transition is a continuous second-order phase transition. The hat notation in disorder potential and pairing denotes the normalized matrix, i.e. ${\widehat{\Delta}}=\tau_{1}$.
Figure S2: Diagrammatic representation of (a) the first-order Born approximation and (b) the particle–particle susceptibility.
C.3 C. $t_{\perp}=0$ case
We begin with the simple case $t_{\perp}=0$, where the two bands are degenerate with dispersion $\xi_{\bm{k}}$. In the limit $\Delta_{0}\rightarrow 0$, the normal and anomalous Green’s functions simplify to
@@
Substituting these expressions into Eqs. S23 and S24, we obtain
$\Sigma_{\omega}(\bm{k},i\omega_{n})$ $=$ $-\frac{i\omega_{n}}{|\omega_{n}|}\pi N(0)n_{\mathrm{imp}}V^{2},$ (S26)
@@
$\Sigma_{\Delta}(\bm{k},i\omega_{n})$ $=$ $\frac{\Delta_{0}}{|\omega_{n}|}\pi N(0)\frac{n_{\mathrm{imp}}V^{2}}{2}\,\mathrm{Tr}\Big[{\widehat{V}}\,{\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big],$ (S27)
where we used $\int_{\bm{k}}=N(0)\int d\xi$, with $N(0)$ the density of states at the Fermi energy.
The trace part in Eq. S27 can be simplified using
$\mathrm{Tr}\Big[{\widehat{V}}\,{\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big]=2-\frac{1}{2}\mathrm{Tr}\Big[{\widehat{F}}_{c}^{\dagger}{\widehat{F}}_{c}\Big],$ (S28)
where the superconducting fitness function is defined as
@@
Substituting Eq. S28 into Eq. S27, we obtain
$\Sigma_{\Delta}(\bm{k},i\omega_{n})=\frac{\Delta_{0}}{|\omega_{n}|}\pi N(0)n_{\mathrm{imp}}V^{2}\left[1-\frac{1}{4}\mathrm{Tr}\Big({\widehat{F}}_{c}^{\dagger}{\widehat{F}}_{c}\Big)\right].$ (S30)
The renormalized Matsubara frequency and pairing amplitude are therefore given by
$\tilde{\omega}_{n}$ $=$ $\omega_{n}+\frac{\tilde{\omega}_{n}}{|\tilde{\omega}_{n}|}\Gamma,$ (S31)
$\tilde{\Delta}_{0}$ $=$ $\Delta_{0}+\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|}\left[1-\frac{1}{4}\mathrm{Tr}\Big(\tilde{F}_{c}^{\dagger}\tilde{F}_{c}\Big)\right]\Gamma,$ (S32)
where scattering frequency, $\Gamma\equiv\pi N(0)n_{\mathrm{imp}}V^{2}=\tau^{-1}$ is used.
As we show in the next subsection, the ratio between $\tilde{\omega}_{n}/\tilde{\Delta}_{0}\equiv u_{n}$ is central quantity in determining the disorder-induced suppression of the critical temperature, as obtained from the linearized gap equation. From Eqs. S31S32, one can find
$\frac{\omega_{n}}{\Delta_{0}}=u_{n}\left[1-\frac{\alpha_{\mathrm{dis}}}{\Delta_{0}}\frac{\Gamma}{|u_{n}|}\right]$ (S33)
with defining depairing coefficient
$\alpha_{\mathrm{dis}}=\frac{1}{4}\mathrm{Tr}\Big(\tilde{F}_{c}^{\dagger}\tilde{F}_{c}\Big).$ (S34)
If the condition $\tilde{\omega}_{n}/\tilde{\Delta}_{0}=\omega_{n}/\Delta_{0}$ is satisfied, meaning that the frequency and pairing renormalizations are identical (equivalently, $\alpha_{\mathrm{dis}}=0$), the superconducting state is robust against the corresponding disorder channel.
Finally, for the disorder channels ${\widehat{V}}=\tau_{i}$ defined in Eq. S20 and interlayer pairing ${\widehat{\Delta}}=\tau_{1}$, we find
@@
consequently, $\alpha_{\mathrm{dis}}=0$ for all channels except $\tau_{3}$, while $\alpha_{\mathrm{dis}}=2$ for $\tau_{3}$. This result directly leads to Table 1. The $\tau_{3}$ disorder channel breaks the composite $\mathcal{M}\mathcal{T}$ symmetry, while all other channels preserve it.
C.3.1 Renormalized gap equation and $T_{c}$
@@
where $g>0$ denotes the attractive pairing interaction. The gap equation is achieved by the second order Ginzburg Landau diagram as illustrated in Fig. S2(b).
Replacing the momentum summation by an integral over energies near the Fermi surface and using Eqs. S31-S32, the gap equation can be rewritten as
$\frac{1}{2g}\simeq\pi N(0)T_{c,0}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{1}{|\omega_{n}|}\;\rightarrow\;\frac{\Delta_{0}}{2g}\simeq\pi N(0)T_{c}\!\!\sum_{|\omega_{n}|<\Lambda_{UV}}\!\!\frac{\tilde{\Delta}_{0}}{|\tilde{\omega}_{n}|},$ (S37)
where $\Lambda_{UV}$ is an ultraviolet energy cutoff imposed on the Matsubara frequencies.
Substituting the renormalized parameters from Eq. S33, we obtain
$\frac{1}{2g}\simeq\pi N(0)T_{c}\sum_{|\omega_{n}|<\Lambda_{UV}}\Bigg[\frac{1}{\omega_{n}}-\frac{\alpha_{\mathrm{dis}}}{\omega_{n}^{2}}\Gamma\Bigg]$ (S38)
To perform the Matsubara summation in Eq. S38, we make use of the digamma function $\Psi(z)$ and its derivative $\Psi^{(1)}(z)=d\Psi(z)/dz$ defined as
@@
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}}$ $=$ $-\Bigg[\Psi\!\left(\frac{1}{2}\right)-\Psi\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$ (S39)
@@
$2\pi T_{c}\sum_{\omega_{n}=0}^{\Lambda_{UV}}\frac{1}{\omega_{n}^{2}}$ $=$ $\frac{1}{2\pi T_{c}}\Bigg[\Psi^{(1)}\!\left(\frac{1}{2}\right)-\Psi^{(1)}\!\left(\frac{1}{2}+\frac{\Lambda_{UV}}{2\pi T_{c}}\right)\Bigg]$ (S40)
@@
$\log\left(\frac{T_{c}}{T_{c,0}}\right)$ $\simeq$ $-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma,$ (S41)
@@
Figure S3: Transition temperature $T_{c}$ as a function of the scattering rate $\Gamma$, obtained from Eq. S51. The $\tau_{3}$ disorder channel, which breaks mirror symmetry, suppresses superconductivity (i.e. $\alpha_{\mathrm{dis}}=2$) irrespective of whether $t_{\perp}=0$ or finite.
C.4 D. $t_{\perp}\neq 0$ case
For the case where $t_{\perp}\neq 0$, the band splitting modifies the dispersion to $\xi_{\pm}(\bm{k})=\xi(\bm{k})\pm|t_{\perp}|$, which may modify the expression for the depairing coefficient $\alpha_{\mathrm{dis}}$. Following the framework introduced in Sec. III A, we evaluate the normal and anomalous Green’s functions, $G$ and $F$, in the limit $\Delta_{0}\rightarrow 0$ to calculate the self-energies.
In the band basis, the normal Green’s function is expressed as,
$G_{N}(\bm{k},i\omega_{n})$ $=$ $\frac{1}{2}\left[(i\omega_{n}-\xi_{+}(\bm{k}))^{-1}+(i\omega_{n}-\xi_{-}(\bm{k}))^{-1}\right]\tilde{\tau}_{0}$ (S42)
$+$ $\frac{1}{2}\left[(i\omega_{n}-\xi_{+}(\bm{k}))^{-1}-(i\omega_{n}-\xi_{-}(\bm{k}))^{-1}\right]\tilde{\tau}_{3},$
@@
where the tilde notation denotes Pauli matrices defined in the band basis. The anomalous Green’s function in the same limit is given by,
$\lim_{\Delta_{0}\rightarrow 0}\frac{F(\bm{k},i\omega_{n})}{\Delta_{0}}$ $=$ $G_{N}{\widehat{\Delta}}G_{N}^{*}=G_{N}G_{N}^{*}{\widehat{\Delta}}$, (S43)
@@
where we have used the fact that $G_{N},{\widehat{\Delta}}]=0$, as both operators are diagonal in the band basis (i.e., expressed in $\tilde{\tau}_{0}$ and $\tilde{\tau}_{3}$).
Substituting these expressions into the self-energy equations, we obtain the results for $t_{\perp}\neq 0$. For the frequency self-energy, we find,
$\lim_{\Delta_{0}\rightarrow 0}\mathrm{Tr}\Big[{\widehat{V}}G(\bm{k},i\omega_{n}){\widehat{V}}\Big$] $=$ $\mathrm{Tr}\Big[\left(G_{N}{\widehat{V}}+[{\widehat{V}},G_{N}]\right){\widehat{V}}\Big]=\mathrm{Tr}\Big[G_{N}{\widehat{V}}{\widehat{V}}\Big$] (S44)
$=$ $-\left(\frac{i\omega_{n}+\xi_{+}}{\omega_{n}^{2}+\xi_{+}^{2}}+\frac{i\omega_{n}+\xi_{-}}{\omega_{n}^{2}+\xi_{-}^{2}}\right),$ (S45)
and for the pairing self-energy,
$\lim_{\Delta_{0}\rightarrow 0}\frac{1}{\Delta_{0}}\mathrm{Tr}\Big[{\widehat{V}}F(\bm{k},i\omega_{n}){\widehat{V}}^{*}{\widehat{\Delta}}^{\dagger}\Big]$ $=$ $\mathrm{Tr}\Big[\big([{\widehat{V}},G_{N}G_{N}^{*}]+G_{N}G_{N}^{*}{\widehat{V}}\big){\widehat{\Delta}}{\widehat{V}}^{*}{\widehat{\Delta}}^{\dagger}\Big]$ (S46)
$=$ $\mathrm{Tr}\Big[G_{N}G_{N}^{*}{\widehat{V}}{\widehat{\Delta}}{\widehat{V}}^{*}{\widehat{\Delta}}^{\dagger}\Big]$